# Ensemble noise properties of the European Pulsar Timing Array

### v2

@@ Abstract
The null hypothesis in Pulsar Timing Array (PTA) analyses includes assumptions about ensemble properties of pulsar time-correlated noise. These properties are encoded in prior probabilities for the amplitude and the spectral index of the power-law power spectral density of temporal correlations of the noise. In this work, we introduce a new procedure for numerical marginalisation over the uncertainties in pulsar noise priors. The procedure may be used in searches for nanohertz gravitational waves and other PTA analyses to resolve prior misspecification at negligible computational cost. Furthermore, we infer the distribution of amplitudes and spectral indices of the power spectral density of spin noise and dispersion measure variation noise based on the observation of 25 millisecond pulsars by the European Pulsar Timing Array (EPTA). Our results may be used for the simulation of realistic noise in PTAs.
@@
pubyear: 2024pagerange: Ensemble noise properties of the European Pulsar Timing ArrayA
@@ 1 Introduction
Pulsar Timing Arrays (PTAs, Foster & Backer, 1990) are experiments that monitor pulse arrival times from galactic millisecond pulsars with a primary goal of detecting nanohertz-frequency gravitational waves (Sazhin, 1978; Detweiler, 1979). The most promising source of such gravitational waves is the stochastic superposition of inspiralling supermassive binary black holes in the nearby universe (Rosado et al., 2015). Thanks to spacetime metric perturbations from gravitational waves, pulse arrival times experience delays and advances (henceforth, delays). The power spectral density (PSD) of delays $P(f)$ induced by stochastic gravitational waves manifests temporal correlations and corresponds to the background’s characteristic strain spectrum $h_{\text{c}}(f)$. The Fourier frequency of timing delays $f$ is exactly the gravitational wave frequency. For the isotropic stochastic gravitational wave background from circular binaries where the inspiral is driven by gravitational wave emission alone, $h_{\text{c}}\propto f^{-2/3}$ corresponding to $P(f)\propto f^{-13/3}$, highlighting signal prominence towards lower frequencies or, equivalently, longer observational time scales (Renzini et al., 2022). At low frequencies, PTAs are limited by time-correlated “red” noise which is also modelled using a power law. The two most ubiquitous sources of red noise in PTA data are the dispersion measure (DM) variation noise (Keith et al., 2013) and the spin noise (Shannon & Cordes, 2010). DM noise features an additional dependence $P(f)\propto\nu^{-2}$, where $\nu$ is a radio frequency. It belongs to a broader class of noise processes called “chromatic”, referring to the dependence of the signal amplitude on radio frequency. Achromatic red noise that is independent of $\nu$ is also called spin noise because it is associated with irregularities in pulsar rotation. In data of a single pulsar, spin noise acts as a gravitational wave background with a broader range of possible spectral indices. A decisive contribution of the stochastic background to PTA data is ultimately determined through the covariance of the signal across pulsar pairs. The covariance follows the Hellings & Downs (1983) function of pulsar angular separation.
@@
Contemporary PTA analyses are performed using Bayesian inference, where a uniform prior probability is assumed for pulsar-specific red noise parameters. Namely, the log-10 amplitude, $\lg A$, and the spectral index, $\gamma$, of noise PSD. Priors are the models of how likely a parameter with a certain value is to be found in a given data realization. A prior choice is sufficient as long as it is consistent with an observation. Uniform priors on PTA noise parameters are designed to be sufficient for single-pulsar noise analyses. However, all pulsars represent different realizations of data with respect to pulsar noise parameters [^1] While remaining a single realization of data with respect to parameters governing $h_{\text{c}}(f)$ of gravitational wave background.. So, for full-PTA analyses with multiple pulsars, a mismatch between our prior on $(\lg A,\gamma)$ and the observed distribution of these parameters may accumulate across pulsars and lead to biased inference. This case is known as prior misspecification.
@@
Implicitly, noise prior misspecification is shown by Hazboun et al. (2020) to bias measurements of the strain amplitude of the gravitational wave background in PTAs. The authors have mitigated the bias by allowing the data to choose whether red noise is present or absent in a pulsar. Furthermore, in simulations of Goncharov et al. (2021b); Zic et al. (2022), pulsar spin noise with a wide range of $(\lg A,\gamma)$ modelled with the standard uniform priors, not representative of ensemble noise properties, yields evidence for a stochastic signal with the same power-law PSD across pulsars (which is not present in the simulated data). Such a signal, under the assumption of uniform noise priors, has been reported in real PTA data as a possible precursor to Hellings-Downs correlations of the gravitational wave background (Goncharov et al., 2021b; NANOGrav Collaboration, 2020; Chen et al., 2021).
@@
Goncharov et al. (2022) regularised incorrect pulsar noise priors by allowing noise amplitudes of the common-spectrum signal of the putative gravitational wave background to vary across pulsars. By showing that this variance is consistent with zero, the authors confirmed the consistency of the signal with the gravitational wave background. Although the model of Goncharov et al. (2022) provides the working solution and has clear use cases, the ultimate solution to the misspecification of noise priors is to parametrise priors for all relevant noise parameters. Both the statement of the problem and the most general solution are clearly outlined in van Haasteren (2024). The distribution of pulsar noise parameters is inferred simultaneously with a search for Hellings-Downs correlations and other red processes with pulsar-to-pulsar covariance. The technique is shown to mitigate systematic errors introduced by incorrect noise priors.
@@ 2.1 Standard PTA data analysis methodology
Here, $\bm{F}$ are the Fourier sine and cosine basis function and $\bm{a}$ are Fourier amplitudes of red processes. Timing model contributions are similarly modelled via the design matrix $\bm{M}$ and the coefficients $\bm{\epsilon}$ represent timing model parameters. Whereas a type of white noise corresponding to timing delays $\bm{j}$ in units [s] that are the same for all ToAs obtained in the same observing epoch (“jitter” noise) are modelled using the basis $\bm{U}$. Vector $\bm{d}(t)$ corresponds to other deterministic signals as a function of pulsar observation time. Let us denote $\bm{b}\equiv(\bm{a},\bm{\epsilon},\bm{j})\in\bm{\theta}$. Bayes theorem is used to obtain a posterior distribution of model parameters $\mathcal{P}(\bm{\theta}|\bm{\delta t})$ from the likelihood and the prior $\pi(\bm{\theta})$:
@@
Hierarchical inference – the approach of parametrising prior distributions – is already a part of the standard PTA data analysis machinery. It is employed to reduce the parameter space and to make it physically-motivated. Instead of Fourier amplitudes $\bm{a}$ (two per frequency, one per sine term of $\bm{F}$ and one per cosine term) it is more convenient to measure power-law parameters $(A,\gamma)$ of the power spectral density of red processes:
@@
Goncharov et al. (2022), $\bm{\Lambda}_{\text{CP}}=(\mu_{\lg A_{\text{CP}}},\sigma_{\lg A_{\text{CP}}})$  $\mathcal{P}(\bm{\Lambda}_{\text{CP}}|\bm{\delta t})$ Yes Via a prior in individual pulsars Modelling of the common-spectrum red process (pulsar-to-pulsar autocorrelation of the GWB)
@@ 2.2 Hierarchical inference of ensemble noise properties
As part of the standard PTA analysis routine described in Section 2.1, priors $\pi(\lg A,\gamma)$ are assumed to be uniform distributions, $\mathcal{U}$. Typically, $\pi(\lg A)=\mathcal{U}(-20,-12)$, $\pi(\gamma)=\mathcal{U}(0,7)$. As we pointed out in the Introduction, observational data hints that this model may be incorrect. The solution is to propose alternative parametrised models $\pi(\bm{\theta}|\bm{\Lambda})$ which now depend on hyperparameters $\bm{\Lambda}$. Therefore, generalising $(\lg A,\gamma)\in\bm{\theta}$, the posterior becomes
@@ 2.2.1 Marginalised posterior of hyperparameters
To obtain $\mathcal{P}(\bm{\Lambda})$ marginalised over $\bm{\theta}$, we propose a solution based on prior reweighting, the application of importance sampling to priors. The calculation is done in two steps, both of which manifest one global fit to data. First, one obtains $\mathcal{P}(\bm{\theta}^{\prime},\bm{\theta}^{\prime\prime}|\bm{\delta t})$ for every pulsar assuming a fixed prior on noise parameters $\pi(\bm{\theta}^{k}_{i}|{\varnothing})$. Here, $\varnothing$ means that hyperparameters are fixed as in the standard EPTA analysis (EPTA Collaboration & InPTA Collaboration, 2023a). Second, one uses the resulting posterior samples and evidence values ${\cal Z}_{\varnothing,i}$ to construct a likelihood marginalised over all parameters except $\bm{\Lambda}$, $\mathcal{L}(\bm{\delta t}|\bm{\Lambda})$. It is shown in Thrane & Talbot (2019) that the likelihood from Equation 7 can be written in the following form:
@@
In the above equation, $N_{\text{psrs}}$ is the number of pulsars in a PTA, $n_{i}$ is the number of posterior samples obtained for a given pulsar, $\theta^{k}_{i}$ is the $k$’th posterior sample for $i$’th pulsar. Ensemble noise properties of the EPTA presented in this work are obtained based on Equation 8. The limitation of this approach is that it is based on expressing the total PTA likelihood as the product of single-pulsar likelihoods. Therefore, the approach is blind to Hellings-Downs correlations or other inter-pulsar correlations in the timing data. Whereas the presence of temporal correlations with the same $(\lg A,\gamma)$ in all pulsars has to be imposed via a prior in individual pulsars. To model a common-spectrum process (NANOGrav Collaboration, 2020; Goncharov et al., 2021b; Chen et al., 2021) associated with a gravitational wave background (NANOGrav Collaboration, 2023a; EPTA Collaboration & InPTA Collaboration, 2023b; Reardon et al., 2023) in our hierarchical likelihood defined by Equation 8, we impose an additional spin-noise-like term with $\gamma=13/3$ in individual pulsar likelihoods. Improved methods for separating the effect of gravitational wave background from pulsar-intrinsic work may be explored in future work.
@@ 2.2.2 Marginalisation over hyperparameters
where $\bm{\Lambda}_{k}$ are $n_{\text{p}}$ samples from the prior. The reader may notice that the resulting posterior in Equation 9 is the product ($\times$) of the standard PTA posterior and the weight factor. A derivation of Equation 9 is provided in the Appendix. Unlike Equation 8, Equation  9 allows to simultaneously model Hellings-Downs and other inter-pulsar correlations in the data while accounting for the uncertainty in pulsar noise priors. It is used more extensively in the companion paper (Goncharov et al., in prep.), where the likelihood is multiplied by the weight factor from Equation 9, and the rest of the analysis is performed in a standard way. Alternatively, the posterior $\mathcal{P}(\bm{\theta}|\bm{\delta t})$ can be reweighted into a $\bm{\Lambda}$-marginalised posterior using rejection sampling.
@@ 3 Results
In this work, we infer distributions for parameters governing pulsar spin noise, $\pi(\lg A_{\text{SN}},\gamma_{\text{SN}})$, and DM noise, $\pi(\lg A_{\text{DM}},\gamma_{\text{DM}})$, for EPTA DR2. EPTA DR2 contains a common-spectrum stochastic process associated with the gravitational wave background and $\gamma$ consistent with $13/3$ (Goncharov et al., in prep.). If we do not account for this noise term, it may appear as a “quasi-common” red noise (as per the terminology of Goncharov et al., 2022). In other words, as a separate cluster in the distribution of spin noise parameters. Because we are interested in knowing the distribution of pulsar-intrinsic parameters, we include an additional red noise term with $\gamma_{\text{c}}=13/3$ to marginalise over a contribution of the common signal. This is done during the analysis of single pulsar noise corresponding to the first step in Section 2.2. It corresponds to modelling common signal in our data via a prior as per Table 1. Because we are agnostic about $\lg A_{\text{c}}$, the approach imposes a suboptimal uncertainty for measuring $(\lg A,\gamma)$ in pulsars where $\gamma$ is consistent with $13/3\approx 4$. Furthermore, to separate contributions of other noise processes such as chromatic noise (including scattering variations) and band- or system-dependent noise, we adopt single-pulsar noise models and optimal numbers of Fourier frequency bins from the EPTA DR2 noise analysis (EPTA Collaboration & InPTA Collaboration, 2023a).
@@ 3.1 Uniform distribution of noise parameters
We start with the case of uniform distributions for noise parameters and we measure hyperparameters that govern uniform prior range. For DM noise, these are $\text{min}(\lg{A}_{\text{DM}})$, $\text{max}(\lg{A}_{\text{DM}})$, $\text{min}(\gamma_{\text{DM}})$, $\text{max}(\gamma_{\text{DM}})$. We apply the same principle to spin noise (SN) parameters. Pulsar-specific measurements of $(\lg{A},\gamma)$ and the inferred boundaries of the uniform distribution for spin noise and DM noise parameters – with measurement uncertainties – are shown in Figure 1. The first important observation is that the data rules out, with high credibility, the standard values of uniform prior boundaries. The data suggests that pulsar noise parameters are distributed in a more narrow range. Posterior density $\mathcal{P}(\lg A,\gamma)$ from pulsars that do not have evidence for the respective noise term according to the EPTA Collaboration & InPTA Collaboration (2023a) is shown in colour. The data from these pulsars has also contributed to the measurement of hyperparameters. The inferred noise distribution model extends as little as possible to include as many identified noise parameters in pulsars as possible. One may, in principle, extend our model to include selection effects associated with a lack of sensitivity for certain $A,\gamma$ (see Appendix E in Thrane & Talbot, 2019). This would enable excess density of $\pi(\lg A|\bm{\Lambda})$ towards low $\lg A$.
@@
Figure 1: The inferred form of the uniform distribution of pulsar red noise parameters. Black solid contours correspond to $1\sigma$ posterior levels of pulsar noise parameters when analysing single pulsar data, provided that the noise term is resolved in a pulsar based on model selection from EPTA Collaboration & InPTA Collaboration (2023a). Colour corresponds to the posterior density of pulsars where noise terms were not resolved. Our hierarchical analysis yields hyperparameters which determine how pulsar noise parameters are distributed. The shaded rectangular areas correspond to $1\sigma$ credible levels for hyperparameters $(\bm{\min},\bm{\max})$ that govern limits of the uniform distribution based on our hierarchical analysis. Hyperparameter values corresponding to the shaded area and the dashed line represent a row $\mathcal{U}(\bm{\theta}|\textbf{min},\textbf{max})$ in Table 2.

### v3

@@ Abstract
The null hypothesis in Pulsar Timing Array (PTA) analyses includes assumptions about ensemble properties of pulsar time-correlated noise. These properties are encoded in prior probabilities for the amplitude and the spectral index of the power-law power spectral density of temporal correlations of the noise. Because multiple realizations of time-correlated noise processes are found in pulsars, these ensemble noise properties could and should be modelled in the full-PTA observations by parameterising the respective prior distributions using the so-called hyperparameters. This approach is known as the hierarchical Bayesian inference. In this work, we introduce a new procedure for numerical marginalisation over hyperparameters. The procedure may be used in searches for nanohertz gravitational waves and other PTA analyses to resolve prior misspecification at negligible computational cost. Furthermore, we infer the distribution of amplitudes and spectral indices of the power spectral density of spin noise and dispersion measure variation noise based on the observation of 25 millisecond pulsars by the European Pulsar Timing Array (EPTA). Our results may be used for the simulation of realistic noise in PTAs.
@@
pubyear: 2025pagerange: Ensemble noise properties of the European Pulsar Timing ArrayB
@@ 1 Introduction
Pulsar Timing Arrays (PTAs, Foster & Backer, 1990) are experiments that monitor pulse arrival times from galactic millisecond pulsars with a primary goal of detecting nanohertz-frequency gravitational waves (Sazhin, 1978; Detweiler, 1979). The most promising source of such gravitational waves is the stochastic superposition of inspiralling supermassive binary black holes in the nearby universe (Rosado et al., 2015). Thanks to spacetime metric perturbations from gravitational waves, pulse arrival times experience delays and advances (henceforth, delays). The power spectral density (PSD) of delays $P(f)$ induced by stochastic gravitational waves manifests temporal correlations and corresponds to the background’s characteristic strain spectrum $h_{\text{c}}(f)$. The Fourier frequency of timing delays $f$ is exactly the gravitational wave frequency. For the isotropic stochastic gravitational wave background from circular binaries where the inspiral is driven by gravitational wave emission alone, $h_{\text{c}}\propto f^{-2/3}$ corresponding to $P(f)\propto f^{-13/3}$, highlighting signal prominence towards lower frequencies or, equivalently, longer observational time scales (Renzini et al., 2022; Phinney, 2001). At low frequencies, PTAs are limited by time-correlated “red” noise which is also modelled using a power law. The two basic sources of red noise in PTA data are the dispersion measure (DM) variation noise (Keith et al., 2013) and the spin noise (Shannon & Cordes, 2010). DM noise features an additional dependence $P(f)\propto\nu^{-2}$, where $\nu$ is a radio frequency. It belongs to a broader class of noise processes called “chromatic”, referring to the dependence of the signal amplitude on radio frequency. Achromatic red noise that is independent of $\nu$ is also called spin noise because it is associated with irregularities in pulsar rotation. A decisive contribution of the stochastic background to PTA data is ultimately determined through the covariance of the signal across pulsar pairs. The covariance follows the Hellings & Downs (1983) function of pulsar angular separation.
@@
Contemporary PTA analyses are performed using Bayesian inference, where a uniform prior probability is assumed for pulsar-specific red noise parameters. Namely, the log-10 amplitude, $\lg A$, and the spectral index, $\gamma$, of noise PSD. Often the term “prior” is used to refer to our knowledge of a parameter before evidence is taken into account. However, more precisely, priors are the models of how likely a parameter with a certain value is to be found in a given data realization. A prior choice is sufficient as long as it is consistent with an observation. Uniform priors on PTA noise parameters are designed to be sufficient for single-pulsar noise analyses. However, all pulsars represent different realizations of data with respect to pulsar noise parameters [^1] While remaining a single realization of data with respect to parameters governing $h_{\text{c}}(f)$ of gravitational wave background.. So, for full-PTA analyses with multiple pulsars, a mismatch between our prior on $(\lg A,\gamma)$ and the observed distribution of these parameters may accumulate across pulsars and lead to biased inference. This case is known as prior misspecification.
@@
Implicitly, noise prior misspecification is shown by Hazboun et al. (2020) to bias measurements of the strain amplitude of the gravitational wave background in PTAs. The authors have mitigated the bias by allowing the data to choose whether red noise is present or absent in a pulsar. Furthermore, in simulations of Goncharov et al. (2021b); Zic et al. (2022), pulsar spin noise with a wide range of $(\lg A,\gamma)$ modelled with the standard uniform priors, not representative of ensemble noise properties, yields evidence for a stochastic signal with the same power-law PSD across pulsars (which is not present in the simulated data). Such a common-spectrum process (CP), under the assumption of uniform noise priors, has been reported in real PTA data as a possible precursor to Hellings-Downs correlations of the gravitational wave background (Goncharov et al., 2021b; NANOGrav Collaboration, 2020; Chen et al., 2021).
@@
Goncharov et al. (2022) regularised incorrect pulsar noise priors by allowing noise amplitudes of the CP of the putative gravitational wave background to vary across pulsars. By showing that this variance is consistent with zero, the authors confirmed the consistency of the signal with the gravitational wave background. Although the model of Goncharov et al. (2022) provides the working solution and has clear use cases, the ultimate solution to the misspecification of noise priors is to parameterise priors for all relevant noise parameters. Both the statement of the problem and the most general solution are clearly outlined in van Haasteren (2024). The distribution of pulsar noise parameters is inferred simultaneously with a search for Hellings-Downs correlations and other red processes with pulsar-to-pulsar covariance. The technique is shown to mitigate systematic errors introduced by incorrect noise priors.
@@ 2.1 Standard PTA data analysis methodology
Here, $\bm{F}$ are the Fourier sine and cosine basis function and $\bm{a}$ are Fourier amplitudes of red processes. Timing model contributions are similarly modelled via the design matrix $\bm{M}$ and the coefficients $\bm{\epsilon}$ represent timing model parameters. Whereas a type of white noise corresponding to timing delays $\bm{j}$ in units [s] that are the same for all ToAs obtained in the same observing epoch (“jitter” noise) are modelled using the basis $\bm{U}$. Vector $\bm{d}(t)$ corresponds to other deterministic signals as a function of pulsar observation time. Deterministic signals include exponential dips associated with pulse shape changes (Goncharov et al., 2021a). Let us denote $\bm{b}\equiv(\bm{a},\bm{\epsilon},\bm{j})\in\bm{\theta}$. Bayes theorem is used to obtain a posterior distribution of model parameters $\mathcal{P}(\bm{\theta}|\bm{\delta t})$ from the likelihood and the prior $\pi(\bm{\theta})$:
@@
Hierarchical inference – the approach of parameterising prior distributions – is already a part of the standard PTA data analysis machinery. It is employed to reduce the parameter space and to make it physically-motivated. Instead of Fourier amplitudes $\bm{a}$ (two per frequency, one per sine term of $\bm{F}$ and one per cosine term) it is more convenient to measure power-law parameters $(A,\gamma)$ of the power spectral density of red processes:
@@
Goncharov et al. (2022), $\bm{\Lambda}_{\text{CP}}=(\mu_{\lg A_{\text{CP}}},\sigma_{\lg A_{\text{CP}}})$  $\mathcal{P}(\bm{\Lambda}_{\text{CP}}|\bm{\delta t})$ Yes Via a prior in individual pulsars Modelling of the common-spectrum process (CP, pulsar-to-pulsar autocorrelation of the GWB)
@@ 2.2 Hierarchical inference of ensemble noise properties
As part of the standard PTA analysis routine described in Section 2.1, priors $\pi(\lg A,\gamma)$ are assumed to be uniform distributions, $\mathcal{U}$. For example, $\pi(\lg A)=\mathcal{U}(-20,-12)$, $\pi(\gamma)=\mathcal{U}(0,7)$. As we pointed out in the Introduction, previous studies suggest that this model may be incorrect. The solution is to propose alternative parameterised models $\pi(\bm{\theta}|\bm{\Lambda})$ which now depend on hyperparameters $\bm{\Lambda}$. Therefore, generalising $(\lg A,\gamma)\in\bm{\theta}$, the posterior becomes
@@ 2.2.1 Marginalised posterior of hyperparameters
To obtain $\mathcal{P}(\bm{\Lambda})$ marginalised over $\bm{\theta}$, we propose a solution based on prior reweighting, the application of importance sampling to priors. The calculation is done in two steps, both of which manifest one global fit to data. First, one obtains $\mathcal{P}(\bm{\theta}^{\prime},\bm{\theta}^{\prime\prime}|\bm{\delta t})$ for every pulsar assuming a fixed prior on noise parameters $\pi(\bm{\theta}^{k}_{i}|{\varnothing})$. It is referred to as the proposal distribution. Here, $\varnothing$ means that hyperparameters are fixed as in the standard EPTA analysis (EPTA Collaboration & InPTA Collaboration, 2023a). Second, one uses the resulting posterior samples and evidence values ${\cal Z}_{\varnothing,i}$ to construct a likelihood marginalised over all parameters except $\bm{\Lambda}$, $\mathcal{L}(\bm{\delta t}|\bm{\Lambda})$. It is shown in Thrane & Talbot (2019) that the likelihood from Equation 7 can be written in the following form:
@@
In the above equation, $N_{\text{psrs}}$ is the number of pulsars in a PTA, $n_{i}$ is the number of posterior samples obtained for a given pulsar, $\theta^{k}_{i}$ is the $k$’th posterior sample for $i$’th pulsar. Ensemble noise properties of the EPTA presented in this work are obtained based on Equation 8. The limitation of this approach is that it is based on expressing the total PTA likelihood as the product of single-pulsar likelihoods. Therefore, the approach is blind to Hellings-Downs correlations or other inter-pulsar (angular) correlations in the timing data. Whereas the presence of temporal correlations with the same $(\lg A,\gamma)$ in all pulsars has to be imposed via a prior in individual pulsars. To model CP (NANOGrav Collaboration, 2020; Goncharov et al., 2021b; Chen et al., 2021) associated with the GWB (NANOGrav Collaboration, 2023a; EPTA Collaboration & InPTA Collaboration, 2023b; Reardon et al., 2023) in our hierarchical likelihood defined by Equation 8, we impose an additional spin-noise-like term with $\gamma=13/3$ in individual pulsar likelihoods. The methodology and implications of this choice are discussed in Appendix A. Improved methods for separating the effect of gravitational wave background from pulsar-intrinsic work may be explored in future work.
@@ 2.2.2 Marginalisation over hyperparameters
where $\bm{\Lambda}_{k}$ are $n_{\text{p}}$ samples from the prior. The reader may notice that the resulting posterior in Equation 9 is the product ($\times$) of the standard PTA posterior and the weight factor. A derivation of Equation 9 is provided in the Appendix. Unlike Equation 8, Equation  9 allows to simultaneously model Hellings-Downs and other inter-pulsar correlations in the data while accounting for the uncertainty in pulsar noise priors. It is used more extensively in the companion paper (Goncharov et al., 2024), where the likelihood is multiplied by the weight factor from Equation 9, and the rest of the analysis is performed in a standard way. Alternatively, the posterior $\mathcal{P}(\bm{\theta}|\bm{\delta t})$ can be reweighted into a $\bm{\Lambda}$-marginalised posterior using rejection sampling.
@@ 3 Results
In this work, we infer distributions for parameters governing pulsar spin noise, $\pi(\lg A_{\text{SN}},\gamma_{\text{SN}})$, and DM noise, $\pi(\lg A_{\text{DM}},\gamma_{\text{DM}})$, for EPTA DR2. EPTA DR2 contains CP associated with the GWB, such that $\gamma_{\text{CP}}$ is consistent with $13/3$ (Goncharov et al., 2024). If we do not model this term, it may appear as a “quasi-common” red noise (as per the terminology of Goncharov et al., 2022). In other words, as a separate cluster in the distribution of spin noise parameters. Because we are interested in knowing the distribution of pulsar-intrinsic parameters, we include an additional red noise term with $\gamma_{\text{c}}=13/3$ to marginalise over a contribution of the common signal. This is done during the analysis of single pulsar noise corresponding to the first step in Section 2.2. It corresponds to modelling common signal in our data via a prior as per Table 1. Because we are agnostic about $\lg A_{\text{CP}}$, the approach imposes a suboptimal uncertainty for measuring $(\lg A,\gamma)$ in pulsars where $\gamma$ is consistent with $13/3\approx 4$. Furthermore, to separate contributions of other noise processes such as chromatic noise (including scattering variations) and band- or system-dependent noise, we adopt single-pulsar noise models and optimal numbers of Fourier frequency bins from the EPTA DR2 noise analysis (EPTA Collaboration & InPTA Collaboration, 2023a).
@@ 3.1 Uniform distribution of noise parameters
We start with the case of uniform distributions for noise parameters and we measure hyperparameters that govern uniform prior range. For DM noise, these are $\text{min}(\lg{A}_{\text{DM}})$, $\text{max}(\lg{A}_{\text{DM}})$, $\text{min}(\gamma_{\text{DM}})$, $\text{max}(\gamma_{\text{DM}})$. We apply the same principle to spin noise (SN) parameters. Pulsar-specific measurements of $(\lg{A},\gamma)$ and the inferred boundaries of the uniform distribution for spin noise and DM noise parameters – with measurement uncertainties – are shown in Figure 1. The first important observation is that the data rules out, with high credibility, the standard values of uniform prior boundaries. The data suggests that pulsar noise parameters are distributed in a more narrow range. Posterior density $\mathcal{P}(\lg A,\gamma)$ from pulsars that do not have evidence for the respective noise term according to the EPTA Collaboration & InPTA Collaboration (2023a) is shown in colour. The data from these pulsars has also contributed to the measurement of hyperparameters.
@@
Figure 1: The inferred form of the uniform distribution of pulsar red noise parameters. Black solid contours correspond to $1\sigma$ posterior levels of pulsar noise parameters when analysing single pulsar data, provided that the noise term is resolved in a pulsar based on model selection from EPTA Collaboration & InPTA Collaboration (2023a) (where inference is performed without the CP term, therefore our solid contours for SN may be different to those in the reference). Colour corresponds to the posterior density of pulsars where noise terms were not resolved. Our hierarchical analysis yields hyperparameters which determine how pulsar noise parameters are distributed. The shaded rectangular areas correspond to $1\sigma$ credible levels for hyperparameters $(\bm{\min},\bm{\max})$ that govern limits of the uniform distribution based on our hierarchical analysis. Hyperparameter values corresponding to the shaded area and the dashed line represent a row $\mathcal{U}(\bm{\theta}|\textbf{min},\textbf{max})$ in Table 2.